OpenStudy (anonymous):

simplify ln e ^7 ?????????

hartnn (hartnn):

use the property that , \(\huge \ln x^n =n\ln x\) and ln e =1.

OpenStudy (anonymous):

it looks like this... I think its 7 right? 

OpenStudy (anonymous):

@hartnn the "e" is lowered

hartnn (hartnn):

i think its  then 7 is correct,